Question 17 3 pts Small Sample Mean Problem. Each year the Heritage Foundation creates an Index of Economic Freedom for countIs this the right answer?
Answer #1

Solution :

Your answer is right .

This is the right tailed test .

sample size = n = 31

Degrees of freedom = df = n - 1 = 30

t\alpha /2,df = t 0.10,30 = 1.310